OpenStudy (anonymous):

If A is invertible, then Adj(A) is invertible and [Adj(A)]^-1 = (1/det(A)) A = adj(A^-1)

OpenStudy (anonymous):

@Lyrae help me

OpenStudy (anonymous):

If A is invertible, then Adj(A) is invertible and \[[Adj(A)]^{-1} = \frac{ 1 }{ \det(A) } A = adj (A^{-1})\]

OpenStudy (anonymous):

prove it

OpenStudy (mathmath333):

take small 2*2 matrix for ex this one \(\large\tt \color{black}{\left[\begin{matrix} 2 & 2 \\ 3 & 2 \end{matrix}\right]}\)

OpenStudy (zarkon):

See ... The Laplace expansion

OpenStudy (anonymous):

i'm confused :(

OpenStudy (anonymous):

\[A^{-1} = \frac{ 1 }{ \det(A) }A \times adj (A)\] \[A A^{-1} = \frac{ 1 }{ \det(A) } A \times adj(A)\] \[I = \frac{ 1 }{ \det(A) } A \times A adj(A)\] \[I [adj(A)]^{-1} = \frac{ 1 }{ \det(A) } A\] So \[[adj(A)]^{-1} = \frac{ 1 }{ \det(A) } A\]

OpenStudy (zarkon):

What you have typed is not correct (small errors)

OpenStudy (zarkon):

look at the first line again

OpenStudy (anonymous):

ahhh the first line would be \[A^{-1} = \frac{ 1 }{ \det(A) } adj(A)\]

OpenStudy (anonymous):

just typo :)

OpenStudy (lyrae):

\[A \space adj(A) = \det(A)\space I\]Proofs for this property can be fount on the internet http://thejuniverse.org/PUBLIC/LinearAlgebra/LOLA/laplace/adjoint.html for example. From this we can derive If and only if A is invertible \[A^{-1}A \space adj(A) = adj(A) = \det(A) \space A^{-1}\]\[\frac{ 1 }{ adj(A) } = adj(A)^{-1} = \frac{ 1 }{\det(A) A^{-1}} = \frac{ 1 }{ \det(A) } A\] Similarly (from the same property) \[A^{-1}adj(A^{-1}) =\det(A^{-1}) \space I = \frac{ 1 }{ det(A) } I\]\[AA^{-1}adj(A^{-1}) = adj(A^{-1}) = \frac{ 1 }{ \det(A) } \space A\] thus \[adj(A)^{-1} =\frac{ 1 }{ \det(A) } = adj(A^{-1})\]

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

which is given by the The Laplace expansion

OpenStudy (zarkon):

now replace \(A^{-1}\) by \(A\) then \[A^{-1} = \frac{ 1 }{ \det(A) } adj(A)\] becomes \[A = \frac{ 1 }{ \det(A^{-1}) } adj(A^{-1})\] and then \[adj(A^{-1})=det(A^{-1})A\] \[adj(A^{-1})=\frac{1}{det(A)}A\]